# Faithful Logic Embeddings in HOL – A recipe to have it all: deep and shallow, automated and interactive, heavy and light, proofs and counterexamples, meta and object level

### v1

@@ 10 Conclusion
This paper has shown how interlinked deep and (varying degrees of) shallow logic embeddings can be modeled and implemented in classical higher-order logic, as provided by many modern proof-assistant systems. This technique allows for the mechanization of faithfulness proofs between these embeddings, and, as has been shown in this paper, can even support the automation of such proofs.
The presented technique is interesting for future research in several ways, and is particularly suited to support logic education due to its conciseness, elegance, and support for interactive and automated theorem proving and counterexample finding. The latter was a motivation to illustrate the approach in this paper using simple propositional modal logic, which is routinely used as introductory material in logic courses worldwide. The provided Isabelle/HOL sources should be easily reusable in such a context, and only some basic features of Isabelle/HOL have been used to additionally support this idea.
@@ 10.0.1 Acknowledgements
Thanks are due to all the LogiKEycontributors over the past ten to fifteen years.
Appendix 0.A Preliminaries
Figure 7 presents preliminaries that are uniform for all embeddings studied and which are imported by all Isabelle/HOL theory files presented in this paper.
Two HOL base types w and $\mathcal{S}$ are declared (in lines 4-5) to denote a (nonempty) set of possible worlds and a (nonempty) set of propositional constant symbols, the signature common to all embeddings discussed in this paper. Then (in line 6) the exemplary signature symbols p, p and r are introduced, and type synonyms are declared (in lines 7-8): $\mathcal{W}$ for sets of possible worlds (universe), $\mathcal{R}$ for binary relations on $W$ (accessibility relations), and $\mathcal{V}$ for evaluation functions that assign truth values to the symbols $a\in\mathcal{S}$ for each world $w$. Then some well-known relational properties are introduced as abbreviations (in lines 11-19), and the guarded quantifier notation $\forall w{:}W.\,\varphi$ is defined to stand for $\forall w.\,W\,w\longrightarrow\varphi$ ($W$ is of type $\mathcal{W}$). Finally, the useful backward implication $\longleftarrow$ is defined, which is unfortunately missing in Isabelle/HOL.
Figure 7: Preliminaries.
Appendix 0.B Alternative experiments

### v2

@@ 10 Conclusion
This paper has demonstrated how the simultaneous deployment of deep and shallow embeddings of various degrees in classical higher-order logic can be achieved. This technique allows for the mechanization of faithfulness proofs between the interlinked embeddings, and, as has been shown in this paper, this technique can even support the automation of such proofs.
The work presented is relevant to future research in various areas. Due to its conciseness and elegance, as well as its support for interactive and automated theorem proving and counterexample finding, it can also be used to support logic education effectively. The latter application direction was the main motivation for illustrating this approach using simple propositional modal logic in this paper. This logic is routinely used as introductory material in logic courses worldwide. The Isabelle/HOL sources provided and explained in this paper should be easily reusable to support such courses, and they should also be easily transferable to other higher-order proof assistant systems.
@@ 10.0.1 Acknowledgements
I am grateful to the reviewers of this paper for their valuable feedback. I would also like to thank D. Fuenmayor, D. Kirchner, L. Pasetto and E. Zalta for their comments and all those who have contributed to LogiKEy in the past. My thanks also go to Deepl for the fruitful exchange that (hopefully) helped to improve the quality of the wording of this paper.
Appendix 0.A Alternative experiments